(D3 – 1)y = 3cos2x


1
Expert's answer
2021-06-11T10:11:44-0400
"y'''-y=3\cos(2x)"

Homogeneous Equation


"y'''-y=0"

The characteristic (auxiliary) equation


"r^3-1=0"

"(r-1)(r^2+r+1)=0"

"r_1=1, r_2,3=-\dfrac{1}{2}\pm i\dfrac{\sqrt{3}}{2}"

"y_h=c_1e^{x}+e^{-x/2}\bigg(c_2\cos(\dfrac{\sqrt{3}}{2}x)+c_3\sin(\dfrac{\sqrt{3}}{2}x)\bigg)"

"y_p=A\sin(2x)+B\cos(2x)"

"y_p'=2A\cos(2x)-2B\sin(2x)"

"y_p''=-4A\sin(2x)-4B\cos(2x)"




"y_p'''=-8A\cos(2x)+8B\sin(2x)"

Then


"-8A\cos(2x)+8B\sin(2x)"

"-A\sin(2x)-B\cos(2x)=3\cos(2x)"

"8B-A=0"

"-8A-B=3"

"A=-\dfrac{24}{65}"

"B=-\dfrac{3}{65}"

Therefore


"y(x)=c_1e^{x}+e^{-x/2}\bigg(c_2\cos(\dfrac{\sqrt{3}}{2}x)+c_3\sin(\dfrac{\sqrt{3}}{2}x)\bigg)"

"-\dfrac{24}{65}\sin(2x)-\dfrac{8}{65}\cos(2x)"